OpenStudy (anonymous):

Subtract 4x/x+1-3x+5/x+1

OpenStudy (anonymous):

can you write the formula out for me?

OpenStudy (anonymous):

@Knivez \[\frac{ 4x }{ x+1}-\frac{ 3x+5 }{ x+1 }\]

jimthompson5910 (jim_thompson5910):

The denominators are the same. So you can subtract the numerators over the common denominator like so \[\Large \frac{ 4x }{ x+1}-\frac{ 3x+5 }{ x+1 }\] \[\Large \frac{ 4x -(3x+5) }{ x+1 }\] \[\Large \frac{ 4x -3x-5 }{ x+1 }\] \[\Large \frac{ 1x-5 }{ x+1 }\] \[\Large \frac{ x-5 }{ x+1 }\]